# Byzantine Distributed Source Coding

### v1

@@ Abstract
We study the distributed source coding problem with $k$ users of which at most $s$ may be controlled by an adversary and characterize the set of functions of the sources the decoder can reconstruct robustly in the following sense – if the users behave honestly, the function is recovered with high probability (w.h.p.); if they behave adversarially, w.h.p, either one of the adversarial users will be identified or the function is recovered with vanishingly small distortion.
@@ I Introduction
In the distributed source coding problem, spatially distributed users observe correlated sources and send encodings of their observations to a central decoder which, using these messages and any source observation (side-information) of its own, attempts to reconstruct all the sources or a function of them. Following the seminal work of Slepian and Wolf [1], a wide variety of settings has been studied (e.g., see [2] and references therein). The focus of these works is mainly on understanding the amount of communication required, i.e., how short can the messages from the users be so that the decoder can reconstruct the sources at the desired fidelity. In this paper we turn our attention from limited communication to limited trust. Specifically, the decoder can no longer trust all the users to act faithfully. Out of $k$ users, up to $s\leq k$ may be controlled by an adversary (we call these users malicious); the upper bound $s$ is assumed to be known beforehand, but the identity of which users are controlled by the adversary is unknown to the decoder and the honest users. We ask what, if anything, can the central server hope to learn about the sources in such a setting.
In this paper we are interested in the following form of robust recovery (see Figure 1): the decoder should either be able to compute a desired function $f$ of the sources or it must be able to identify a malicious user (if any). The main result will be a characterization of all functions $f$ for which this is possible for a given joint distribution of the sources. Note that if a malicious user is identified, that user may be removed to yield a new instance of the problem with $k-1$ users of whom at most $s-1$ are malicious and the process continued.
@@
Figure 1: At most $s$ out of the $k$ users are malicious (i.e., controlled by an adversary). The decoder, with high probability, either identifies a malicious user or outputs a substantially correct estimate of $Z^{n}$, where $Z=f(X_{1},\ldots,X_{k},Y)$. The main result is a characterization of functions $f$ for which this is possible for a given $P_{X_{1},\ldots,X_{k}Y}$. If the decoder identifies a malicious user, that user can be removed to get a new instance of the problem with $k-1$ users of which at most $s-1$ are malicious and the process repeated.
@@
In [4], the above was extended to the case of $k=2$ senders with at most $s=1$ corrupt user using a simple idea – for a joint distribution $P_{X_{1}X_{2}Y}$, suppose the users observe $X_{1}$ and $X_{2}$ (dropping the vector notation for convenience), respectively, and the decoder has side-information $Y$. Using the single-user scheme, the decoder can first robustly recover $X_{1}\searrow Y$ (or determine that user 1 is malicious); then using what it has learned, robustly recover $X_{2}\searrow Y_{1}$ (or determine that user 2 is malicious), where $Y_{1}:=(Y,X_{1}\searrow Y)$; then recover $X_{1}\searrow Y_{2}$ (or determine that user 1 is malicious), where $Y_{2}:=(Y_{1},X_{2}\searrow Y_{1})$ and so on. This process can be shown to saturate in a finite number of steps (which depends on the cardinalities of the sources). Thus, either the decoder recovers this (saturated) function or identifies the adversary. Notice that if the adversary is identified, the decoder can proceed with the knowledge that the other user is honest. A consequence of the results in the present paper is that this function (along with $Y$) is the most a decoder can learn robustly111[4] did not formulate the precise question we formulate here (since the goal there was a signature/byzantine broadcast scheme which imposes further restrictions) and the converses proved there do not directly imply an optimality for the formulation here..
@@
We are unaware of works other than [3, 4] which are directly related to our problem. Broadly related are works on byzantine users in multiterminal information theory and network coding such as [5, 6, 7, 8, 9, 10] apart from the extensive literature in the areas of cryptography and distributed computing.
@@ IV The $(k=2,s=1)$ case
Note that the above lemma does not immediately lead to a proof of achievability of Theorem 1 since (i) the adversary need not use a DMC and (ii) even if a DMC is used, the adversary could select one which induces a distribution too close to $\mathcal{V}_{1}\cap\mathcal{V}_{2}$ for the decoder to detect reliably who the adversary is (note that the adversary may select its attack for each given decoder, i.e., for each $\delta,n$). We employ the method of types [12, Chapter 2] to show the following (in Section VII-B4):
@@
For the sake brevity, here we present a proof of the lemma when the $W_{1}$ is a DMC; the proof for the general case (in Section VII-B4) uses an additional symmetrization trick via a random permutation of $[n]$; the same idea was used in the achievability proof of [4]. Suppose $W_{1}$ is a DMC, i.e., let $W_{1}(\overline{x}_{1}^{n}|x_{1}^{n})=W_{\mathsf{dmc}}^{n}(\overline{x}_{1}^{n}|x_{1}^{n}):=\prod_{t=1}^{n}W_{\mathsf{dmc}}(\overline{x}_{1,t}|x_{1,t})$ for some channel $W_{\mathsf{dmc}}$. Then, part (i) of the lemma follows from standard properties of types ([12, Lemma 2.6] and Pinsker’s inequality) since $(X_{1}^{n},X_{2}^{n},Y^{n},\overline{X}_{1}^{n})\sim P_{X_{1}X_{2}Y}W_{\mathsf{dmc}}$ i.i.d. To show part (ii), we use the following technical lemma which is a continuous version of Lemma 2. Notice that (part (i) of) Lemma 2 says that there is a $g$ such that for every $W_{\overline{X}_{1}|X_{1}}$ for which the induced p.m.f. $Q_{\underline{X}_{1}\underline{X}_{2}\underline{Y}}(\underline{x}_{1},\underline{x}_{2},\underline{y}):=\sum_{x_{1}}P_{X_{1}X_{2}Y}(x_{1},\underline{x}_{2},\underline{y})W_{\overline{X}_{1}|X_{1}}(\underline{x}_{1}|x_{1})$ belongs to $\mathcal{V}_{1}\cap\mathcal{V}_{2}$ (which is another way of saying that there is a corresponding $W_{\overline{X}_{2}|X_{2}}$ which induces the same $Q_{\underline{X}_{1}\underline{X}_{2}\underline{Y}}$), we have $\mathbb{P}(f(X_{1},X_{2},Y)\neq g(\overline{X}_{1},X_{2},Y))=0$ under $P_{X_{1}X_{2}Y}W_{\overline{X}_{1}|X_{1}}$. The following lemma (proved in Section VII-B3) generalizes this:
@@
where (b) is a consequence of the fact that $\mathbb{P}(d_{\tt{TV}}(\mathbf{p}_{X_{1}^{n},X_{2}^{n},Y^{n},\overline{X}_{1}^{n}},P_{X_{1}X_{2}Y}W_{\mathsf{dmc}})>\delta)=2^{-\Omega(n)}$ which follows from [12, Lemma 2.6] and Pinsker’s inequality.
@@ VII-B1 Method of Types
We recall some properties from [12, Chapter 2]. Let $X$ and $Y$ be two jointly distributed random variables according to a joint type $P_{XY}\in\mathcal{P}_{n}(\mathcal{X}\times\mathcal{Y})$. For $(x^{n},y^{n})\in\mathcal{T}^{n}_{XY}$, a distribution $Q$ on $\mathcal{X}$, we have
@@ C-1 Preliminaries
[4, Lemma 4]. The decoder $\mathsf{decode}_{(2,1)}(P_{UVW},u^{n},v^{n},w^{n})$ is defined below: On receiving the reports $u^{n},v^{n}$ from users 1 and 2, respectively, and side-information $w^{n}$,
@@
[4, Lemma 4] If $(U_{t},V_{t},W_{t})$ are sampled i.i.d. from a distribution $P_{UVW}$ for $t\in[n]$, then for any $\delta>0$, there exist parameters $\gamma_{0},\gamma_{1},\dots,\gamma_{|\mathcal{U}||\mathcal{V}|}>0$ such that the following holds:
@@ C-A The$(k=2,s=1)$ case
To show that if $f(X_{1},X_{2},Y)$ is a function of $Y^{\ast}$, then $f$ is $1$-viable, it suffices to show that $Y^{\ast}$ is $1$-viable. This is true since $Y^{\ast}=(Y,X_{1}\searrow Y^{|\mathcal{X}_{1}||\mathcal{X}_{2}|},X_{2}\searrow Y^{|\mathcal{X}_{1}||\mathcal{X}_{2}|})$ is recoverable using Protocol 1, as shown in [4, Lemma 4] and every recoverable function is $1$-viable using Theorem 1. In Lemma 22 below, we show the converse, i.e., that if $f$ is $1$-viable, then, $f$ is a function of $Y^{\ast}$. ∎
@@ C-B The general $s=1$ case:
The decoder receives from users $(\underline{X}^{n}_{[k]},Y^{n})$. The decoder chooses any two users $i,j\in[k]$, assigns $(Y^{n},\underline{X}^{n}_{\{i,j\}^{c}})$ as the side information and runs $\mathsf{decode}_{(2,1)}$ with $P_{U,V,W}=P_{X_{i},X_{j},(Y,X_{\{i,j\}^{c}})}$, $u^{n}=\underline{X}_{i}^{n}$, $v^{n}=\underline{X}_{j}^{n}$. If all users are honest, then $\underline{X}^{n}_{i}=X^{n}_{i}$ for every $i\in[k]$. By part (1) and sub-part (b) of part (2) of Lemma 20, if $\mathsf{decode}_{(2,1)}$ doesn’t terminate, then it outputs $Y^{\ast}_{ij}$ w.h.p. and therefore $\mathsf{decode}_{(k,1)}$ outputs $G^{\ast}=h_{ij}(Y^{\ast}_{ij})$ and terminates. Note that $\mathsf{decode}_{(k,1)}$ outputs $G^{\ast}$ if for any $i,j\in[k]$, $\mathsf{decode}_{(2,1)}$ outputs $Y^{\ast}_{ij}$. Now, we consider the case where, for every $i,j\in[k]$, $\mathsf{decode}_{(2,1)}$ terminates before outputting $Y_{ij}^{\ast}$. Whenever $\mathsf{decode}_{(2,1)}$ terminates before outputting $Y^{\ast}_{ij}$, it means that either the corrupt party is not in $\{i,j\}$ and caused $\mathsf{decode}_{(2,1)}$ to implicate one of $\{i,j\}$ due to corrupt side-information, or, all users providing the side-information are honest and the implicated party is the corrupt one. In either case, the party in $\{i,j\}$ who was not implicated can be exonerated, and therefore can be added to the set $S$, which can be thought of as the set of ’trusted users’. In this way, every time $\mathsf{decode}_{(2,1)}$ doesn’t output $Y^{\ast}_{ij}$, a new party is exonerated and is added to the set $S$. When $|S|=k-2$, the only two parties which could be correct are in $[k]\setminus S$. When the last run of $\mathsf{decode}_{(2,1)}$ also doesn’t output $Y^{\ast}_{ij}$, the user who is blamed is therefore the corrupt party with high probability. ∎

### v2

# Byzantine Distributed Function Computation

@@ Abstract
We study the distributed function computation problem with $k$ users of which at most $s$ may be controlled by an adversary and characterize the set of functions of the sources the decoder can reconstruct robustly in the following sense – if the users behave honestly, the function is recovered with high probability (w.h.p.); if they behave adversarially, w.h.p, either one of the adversarial users will be identified or the function is recovered with vanishingly small distortion.
@@ I Introduction
In the distributed source coding problem, spatially distributed users observe correlated sources and send encodings of their observations to a central decoder which, using these messages and any source observation (side-information) of its own, attempts to reconstruct all the sources or a function of them. Following the seminal work of Slepian and Wolf [1], a wide variety of settings has been studied (e.g., see [2] and references therein). The focus of these works is mainly on understanding the amount of communication required, i.e., how short can the messages from the users be so that the decoder can reconstruct the sources at the desired fidelity. In this paper we turn our attention from limited communication to limited trust. Specifically, the decoder can no longer trust all the users to act faithfully. Out of $k$ users, up to $s\leq k$ may be controlled by an adversary (we call these users malicious); the upper bound $s$ is assumed to be known beforehand, but the identity of which users are controlled by the adversary is unknown to the decoder and the honest users. We ask what, if anything, can the central decoder hope to learn about the sources in such a setting. A similar question with somewhat different motivations from ours was studied by Kosut and Tang [3, 4, 5, 6]. We will discuss the similarities and differences with the problem we study towards the end of this section.
In this paper we are interested in the following form of robust recovery (see Figure 1): the decoder should either be able to compute a desired function $f$ of the sources or it must be able to identify a malicious user (if any). The main result will be a characterization of all functions $f$ for which this is possible for a given joint distribution of the sources. Note that if a malicious user is identified, that user may be removed to yield a new instance of the problem with $k-1$ users of whom at most $s-1$ are malicious and the process continued. We do not impose any communication restrictions on the users. In fact, the users send their observations uncoded to the decoder. Furthermore, we assume that the adversary has no additional side-information other than the observations of users it controls.
@@
Figure 1: At most $s$ out of the $k$ users are malicious (i.e., controlled by an adversary who has access to the observations of the malicious users, but no additional side-information). The decoder, with high probability, either identifies a malicious user or outputs a substantially correct estimate of $Z^{n}$, where $Z=f(X_{1},\ldots,X_{k},Y)$. The main result is a characterization of functions $f$ for which this is possible for a given $P_{X_{1},\ldots,X_{k}Y}$. If the decoder identifies a malicious user, that user can be removed to get a new instance of the problem with $k-1$ users of which at most $s-1$ are malicious and the process repeated.
@@
In [8], the above was extended to the case of $k=2$ senders with at most $s=1$ corrupt user using a simple idea – for a joint distribution $P_{X_{1}X_{2}Y}$, suppose the users observe $X_{1}$ and $X_{2}$ (dropping the vector notation for convenience), respectively, and the decoder has side-information $Y$. Using the single-user scheme, the decoder can first robustly recover $X_{1}\searrow Y$ (or determine that user 1 is malicious); then using what it has learned, robustly recover $X_{2}\searrow Y_{1}$ (or determine that user 2 is malicious), where $Y_{1}:=(Y,X_{1}\searrow Y)$; then recover $X_{1}\searrow Y_{2}$ (or determine that user 1 is malicious), where $Y_{2}:=(Y_{1},X_{2}\searrow Y_{1})$ and so on. This process can be shown to saturate in a finite number of steps (which depends on their joint distribution and is at most the product of cardinalities of the alphabets of $X_{1}$ and $X_{2}$). Thus, either the decoder recovers this (saturated) function or identifies the adversary. Notice that if the adversary is identified, the decoder can proceed with the knowledge that the other user is honest. A consequence of the results in the present paper is that this function (along with $Y$) is the most a decoder can learn robustly111[8] did not formulate the precise question we formulate here (since the goal there was a signature/byzantine broadcast scheme which imposes further restrictions) and the converses proved there do not directly imply an optimality for the formulation here..
@@
Related works
Apart from [7, 8], the closest works to ours are by Kosut and Tong on byzantine distributed source coding [3] and the byzantine CEO problem [4, 5, 6]. Distributed source coding with byzantine users studied in [3] is very similar to our setup except that the decoder outputs reconstructions for the observations of all the users with the objective that the reconstructions must be correct for the observations of the honest users. An obvious way to do this is to require the users to send their observations uncoded (as we do in this paper) and output; but this may not be communication efficient. The paper presents the optimal communication rate-tradeoff region of deterministic and randomized fixed-length codes and the optimal sum-rate for variable length codes. Unlike in our formulation, the adversary is allowed to have additional side-information. The main difference with our problem is that the decoder’s aim is not to correctly compute a function of the observations (or declare an outage by identifying a malicious user), but to produce reconstructions for the observations of all users of which those for honest users’ are correct; the decoder may not be able to identify which of its reconstructions are correct. Our focus in this paper is on the feasibility question of which functions are robustly recoverable and we do not study communication efficiency, whereas the feasibility question is trivial for the formulation in [3] and the paper characterizes the optimal rates of communication. Kosut and Tong also studied the byzantine CEO problem, in which the setup is similar except that the decoder’s (CEO) goal is to estimate an underlying source conditioned on which the observations of the users are conditionally independent. They obtained bounds on the rate-region [4, 6] and the error-exponents [5]. In addition, broadly related are works on byzantine users in multiterminal information theory and network coding such as [9, 10, 11, 12, 13, 14, 15, 16] apart from the extensive literature in the areas of cryptography and distributed computing.
@@ IV The $(k=2,s=1)$ case
Note that the above lemma does not immediately lead to a proof of achievability of Theorem 1 since (i) the adversary need not use a DMC and (ii) even if a DMC is used, the adversary could select one which induces a distribution too close to $\mathcal{V}_{1}\cap\mathcal{V}_{2}$ for the decoder to detect reliably who the adversary is (note that the adversary may select its attack for each given decoder, i.e., for each $\delta,n$). We employ the method of types [18, Chapter 2] to show the following (in Section VII-B4):
@@
For the sake brevity, here we present a proof of the lemma when the $W_{1}$ is a DMC; the proof for the general case (in Section VII-B4) uses an additional symmetrization trick via a random permutation of $[n]$; the same idea was used in the achievability proof of [8]. Suppose $W_{1}$ is a DMC, i.e., let $W_{1}(\overline{x}_{1}^{n}|x_{1}^{n})=W_{\mathsf{dmc}}^{n}(\overline{x}_{1}^{n}|x_{1}^{n}):=\prod_{t=1}^{n}W_{\mathsf{dmc}}(\overline{x}_{1,t}|x_{1,t})$ for some channel $W_{\mathsf{dmc}}$. Then, part (i) of the lemma follows from standard properties of types ([18, Lemma 2.6] and Pinsker’s inequality) since $(X_{1}^{n},X_{2}^{n},Y^{n},\overline{X}_{1}^{n})\sim P_{X_{1}X_{2}Y}W_{\mathsf{dmc}}$ i.i.d. To show part (ii), we use the following technical lemma which is a continuous version of Lemma 2. Notice that (part (i) of) Lemma 2 says that there is a $g$ such that for every $W_{\overline{X}_{1}|X_{1}}$ for which the induced p.m.f. $Q_{\underline{X}_{1}\underline{X}_{2}\underline{Y}}(\underline{x}_{1},\underline{x}_{2},\underline{y}):=\sum_{x_{1}}P_{X_{1}X_{2}Y}(x_{1},\underline{x}_{2},\underline{y})W_{\overline{X}_{1}|X_{1}}(\underline{x}_{1}|x_{1})$ belongs to $\mathcal{V}_{1}\cap\mathcal{V}_{2}$ (which is another way of saying that there is a corresponding $W_{\overline{X}_{2}|X_{2}}$ which induces the same $Q_{\underline{X}_{1}\underline{X}_{2}\underline{Y}}$), we have $\mathbb{P}(f(X_{1},X_{2},Y)\neq g(\overline{X}_{1},X_{2},Y))=0$ under $P_{X_{1}X_{2}Y}W_{\overline{X}_{1}|X_{1}}$. The following lemma (proved in Section VII-B3) generalizes this:
@@
where (b) is a consequence of the fact that $\mathbb{P}(d_{\tt{TV}}(\mathbf{p}_{X_{1}^{n},X_{2}^{n},Y^{n},\overline{X}_{1}^{n}},P_{X_{1}X_{2}Y}W_{\mathsf{dmc}})>\delta)=2^{-\Omega(n)}$ which follows from [18, Lemma 2.6] and Pinsker’s inequality.
@@ VII-B1 Method of Types
We recall some properties from [18, Chapter 2]. Let $X$ and $Y$ be two jointly distributed random variables according to a joint type $P_{XY}\in\mathcal{P}_{n}(\mathcal{X}\times\mathcal{Y})$. For $(x^{n},y^{n})\in\mathcal{T}^{n}_{XY}$, a distribution $Q$ on $\mathcal{X}$, we have
@@ VII-B5 Achievability proof of Theorem 1
Acknowledgment
We thank Aaron Wagner for calling our attention to the works by Kosut and Tong on byzantine distributed source coding [3, 4, 5, 6].
@@ C-1 Preliminaries
[8, Lemma 4]. The decoder $\mathsf{decode}_{(2,1)}(P_{UVW},u^{n},v^{n},w^{n})$ is defined below: On receiving the reports $u^{n},v^{n}$ from users 1 and 2, respectively, and side-information $w^{n}$,
@@
[8, Lemma 4] If $(U_{t},V_{t},W_{t})$ are sampled i.i.d. from a distribution $P_{UVW}$ for $t\in[n]$, then for any $\delta>0$, there exist parameters $\gamma_{0},\gamma_{1},\dots,\gamma_{|\mathcal{U}||\mathcal{V}|}>0$ such that the following holds:
@@ C-A The$(k=2,s=1)$ case
To show that if $f(X_{1},X_{2},Y)$ is a function of $Y^{\ast}$, then $f$ is $1$-viable, it suffices to show that $Y^{\ast}$ is $1$-viable. This is true since $Y^{\ast}=(Y,X_{1}\searrow Y^{|\mathcal{X}_{1}||\mathcal{X}_{2}|},X_{2}\searrow Y^{|\mathcal{X}_{1}||\mathcal{X}_{2}|})$ is recoverable using Protocol 1, as shown in [8, Lemma 4] and every recoverable function is $1$-viable using Theorem 1. In Lemma 22 below, we show the converse, i.e., that if $f$ is $1$-viable, then, $f$ is a function of $Y^{\ast}$. ∎
@@ C-B The general $s=1$ case:
The decoder receives from users $(\underline{X}^{n}_{[k]},Y^{n})$. The decoder chooses any two users $i,j\in[k]$, assigns $(Y^{n},\underline{X}^{n}_{\{i,j\}^{c}})$ as the side information and runs $\mathsf{decode}_{(2,1)}$ with $P_{U,V,W}=P_{X_{i},X_{j},(Y,X_{\{i,j\}^{c}})}$, $u^{n}=\underline{X}_{i}^{n}$, $v^{n}=\underline{X}_{j}^{n}$. If all users are honest, then $\underline{X}^{n}_{i}=X^{n}_{i}$ for every $i\in[k]$. By part (1) and sub-part (b) of part (2) of Lemma 20, if $\mathsf{decode}_{(2,1)}$ doesn’t terminate, then it outputs $Y^{\ast}_{ij}$ w.h.p. and therefore $\mathsf{decode}_{(k,1)}$ outputs $G^{\ast}=h_{ij}(Y^{\ast}_{ij})$ and terminates. Note that $\mathsf{decode}_{(k,1)}$ outputs $G^{\ast}$ if for any $i,j\in[k]$, $\mathsf{decode}_{(2,1)}$ outputs $Y^{\ast}_{ij}$. Now, we consider the case where, for every $i,j\in[k]$, $\mathsf{decode}_{(2,1)}$ terminates before outputting $Y_{ij}^{\ast}$. Whenever $\mathsf{decode}_{(2,1)}$ terminates before outputting $Y^{\ast}_{ij}$, it means that either the corrupt party is not in $\{i,j\}$ and caused $\mathsf{decode}_{(2,1)}$ to implicate one of $\{i,j\}$ due to corrupt side-information, or, all users providing the side-information are honest and the implicated party is the corrupt one. In either case, the party in $\{i,j\}$ who was not implicated can be exonerated, and therefore can be added to the set $S$, which can be thought of as the set of ‘trusted users’. In this way, every time $\mathsf{decode}_{(2,1)}$ doesn’t output $Y^{\ast}_{ij}$, a new party is exonerated and is added to the set $S$. When $|S|=k-2$, the only two parties which could be corrupt are in $[k]\setminus S$. When the last run of $\mathsf{decode}_{(2,1)}$ also doesn’t output $Y^{\ast}_{ij}$, the user who is blamed is therefore the corrupt party with high probability. ∎